# A continuous symmetry breaking measure for finite clusters using Jensen-Shannon divergence

### v1

@@ 1 Introduction
Our statistical symmetry breaking measure (SBM), $\mathscr{S}_{T_{\alpha}}$, uses information theory to quantify symmetry loss in a finite atomic cluster under a transformation $T_{\alpha}$. The cluster is represented by a normalized electron-weighted density function $\mu$. The measure is defined as the Jensen-Shannon (JS) divergence [15] between the transformed density $T_{\alpha}[\mu]$ and the original $\mu$, providing a quantitative assessment of the structure’s deviation from the symmetry element $T_{\alpha}$. The measure can be computed for any operation in any atomic cluster. For an undistorted crystal structure, it returns zero for each crystallographic symmetry operation by design. However, with finite size clusters who’s shape breaks symmetry, unit cell distortions, or atomic displacements, it yields a positive value that indicates the degree of SB with respect to each symmetry operation.
In this paper, we derive this measure and conduct numerical tests to evaluate its behavior for different cases, comparing it with another related SBM based on the Kullback-Leibler (KL). We demonstrate that these measures are useful for investigating factors affecting SB, such as boundary shape, perturbation size, and which atom is distorted. They also help identify which symmetries are most violated or preserved when a cluster experience a specific distortion.
2 KL and JS divergences in symmetry breaking
@@
[…]
@@
$\begin{split}D_{\text{JS}}(P,Q)=&\frac{1}{2}D_{\text{KL}}(P\parallel M)+\frac{1}{2}D_{\text{KL}}(Q\parallel M)\\ =&\frac{1}{2}\int_{\mathcal{X}}p(x)\log\frac{p(x)}{m(x)}+\frac{1}{2}\int_{\mathcal{X}}q(x)\log\frac{q(x)}{m(x)}\,dx,\end{split}$ (3)
@@
3 A symmetry breaking measure for finite clusters
3.1 Finite cluster representation
In our proposed model, we represent a finite cluster of atoms as a normalized electron-weighted atomic density function $\mu(x):x\in\mathbb{R}^{3}\to\mathbb{R}$. The electron density at each atom is assumed to be located at the position of the atomic nucleus. However, due to atomic motions, the probability distribution of the atomic density can be approximated by a three-dimensional Gaussian distribution, commonly known as the Debye-Waller approximation [26], weighted by the number of electrons held by the atom. Specifically, for any atom $k$ where the average location of the nucleus is at $\bm{x}_{k}$, the electron-weighted atomic density can be written as
@@ 3.1 Finite cluster representation
where, $e_{k}$ denotes the number of electrons held by atom $k$, $o_{k}$ is its occupancy factor, and $\bm{U}_{k}\in\mathbb{R}^{3\times 3}$ is the anisotropic atomic displacement tensor (ADT). For simplicity, here we treat this distribution as being isotropic, that is, $\bm{U}_{k}=U_{k}\bm{I}_{3}$ is a diagonal matrix, and all the diagonal entries, $U_{k}$, are equal to the $U_{\text{iso}}$ of atom $k$. Here $U_{k}=1/3\times(\bm{U}_{k,11}+\bm{U}_{k,22}+\bm{U}_{k,33})$ can be interpreted as a mean-square displacement averaged over all the three directions [23].
@@
$\tilde{\mu}=\sum_{k=1}^{N}\tilde{\mu_{k}}=\sum_{k=1}^{N}\tilde{\phi_{k}}\,\mathcal{N}(\bm{x}_{k},U_{k}\bm{I}_{3}),$ (5)
@@
$\mu=\frac{\tilde{\mu}}{\|\tilde{\mu}\|_{L_{1}}}=\frac{\tilde{\mu}}{\sum_{k=1}^{N}\tilde{\phi_{k}}}=\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(\bm{x}_{k},U_{k}\bm{I}_{3}),$ (6)
@@
$\begin{split}T_{\alpha}[\mu]=&T_{\alpha}\left[\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(\bm{x}_{k},U_{k}\bm{I}_{3})\right]\\ =&\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(T_{\alpha}[\bm{x}_{k}],U_{k}\bm{I}_{3}).\end{split}$ (7)
3.2 Symmetry breaking measure
@@
[…]
@@
$\mathscr{S}_{T_{\alpha}}^{JS}[\mu]=\frac{1}{2}\int_{\mathcal{X}}\mu(x)\log\frac{\mu(x)}{m(x)}+\frac{1}{2}\int_{\mathcal{X}}T_{\alpha}[\mu](x)\log\frac{T_{\alpha}[\mu](x)}{m(x)}\,dx,$ (9)
where $m$ is an equal mixture of $\mu$ and $T_{\alpha}[\mu]$. Specifically,
$m=\sum_{k=1}^{N}\frac{\phi_{k}}{2}\,\left[\mathcal{N}(\bm{x}_{k},U_{k}\cdot\bm{I}_{3})+\mathcal{N}(T_{\alpha}[\bm{x}_{k}],U_{k}\cdot\bm{I}_{3})\right].$ (10)
$\mathscr{S}_{T_{\alpha}}^{KL}[\mu]$ and $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ both satisfy the three divergence properties:
@@
  • Self identification: $\mathscr{S}_{T_{\alpha}}[\mu]=0$ only if $T_{\alpha}[\mu]=\mu$.
@@
For the case of the JS-SBM, when using a logarithm with a base of 2 in Eq. 9, it is bounded by one. If the natural logarithm is employed, the upper bound becomes $\sqrt{2}$. $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ reaches its upper bound if and only if $\mu$ and $T_{\alpha}$ are disjoint, indicating that the operation $T_{\alpha}$ is broken completely. For a Gaussian mixture $\mu$, which is consistently non-zero, this upper bound will not be realized. However, if $\bm{x}_{k}$ and all $T_{\alpha}[\bm{x}_{k}]$ are sufficiently distant in terms of their $U_{\text{iso}}$’s, $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ can approach close to its bound.
@@
For a high-symmetry heterostructure, such as an infinite crystal with a Face-Centered Cubic (FCC) lattice of nickel, we refer to a set of symmetry operations denoted as $\mathscr{T}=\{T_{i}\}$, which contains transformations capable of yielding an identical object after carried out. We then introduce some changes to this heterostructure. This results in the object losing some of the symmetry operations from its set (and in principle, possibly gaining some new ones) and having a new set of symmetry operations, $\mathscr{T}^{\prime}=\{T^{\prime}_{i}\}$. Strictly speaking, SB of this structure can be defined as the condition where the SBM $\mathscr{S}_{T_{i}}$ of some symmetry operation $T_{i}\in\mathscr{T}$ result in $\mathscr{S}_{T_{i}}>0$.
3.2.1 Symmetry breaking measure of symmetry operators
@@
Determining the SBM of symmetry operators is not trivial, because the symmetry operators can generate more than one symmetry operation. For example, such as the proper axes of rotation $C_{n}$ can lead to $n-1$ operations, specifically $C_{n},C_{n}^{2},\dots,C_{n}^{n-1}$. $C_{n}^{n}$ is considered the “identity” operation, denoted $E$. For example, in structures with n-fold rotational symmetry $C_{n}$, the SBM for these operations are zero ($\mathscr{S}_{C_{n}}=\mathscr{S}_{C_{n}^{2}}=\dots=\mathscr{S}_{C_{n}^{n-1}}=0$). However, if a perturbation disrupts this n-fold symmetry, the SBM for $C_{n}^{i}$ and $C_{n}^{j}$ may differ for some $i,j\in\{1,\dots,n-1\}$. Therefore, we need a measure to quantify how much a particular symmetry operator, such as n-fold rotation, is broken by a symmetry lowering distortion.
To address this, we introduce symmetry breaking measure of operators, $\mathscr{S}_{\mathscr{O}}$, as the average SBM of all the |$|\{\mathscr{O}^{i}\}|$ operations derived from $\mathscr{O}$.
$\mathscr{S}_{\mathscr{O}}=\frac{1}{|\{\mathscr{O}^{i}\}|}\sum_{\mathscr{O}^{i}}\mathscr{S}_{\mathscr{O}^{i}}.$ (12)
For example $\mathscr{S}_{C_{4}}=1/3\sum_{i=1}^{3}\mathscr{S}_{C_{4}^{i}}$. This measure gives a value of 00 when symmetry is preserved, but ranges between 00 and $1$ for the JS-SBM if symmetry is partially broken.
3.2.2 Symmetry breaking measure of a single-atom system
To gain insights, we will consider a structure with a single-atom in the unit cell with $U_{\text{iso}}$ $=U$ and which becomes displaced by $\bm{d}$. We have
$\mu(\bm{x})=$ $\mathcal{N}(\bm{x};\bm{\mu},U\bm{I}_{3}),\,$ (13)
$T_{\bm{d}}[\mu](\bm{x})=$ $\mathcal{N}(\bm{x};\bm{\mu}+\bm{d},U\bm{I}_{3}).$ (14)
@@ 3.2.2 Symmetry breaking measure of a single-atom system
$\mathscr{S}^{KL}_{T_{\bm{d}}}=D_{\text{KL}}\left(\mu\|T_{\bm{d}}[\mu]\right)=\frac{d^{2}}{2U},$ (15)
where $d=\|\bm{d}\|_{L_{2}}$ is the displacement distance. Details of the derivation can be found in the supplemental materials. This result suggests that for small distortions $\bm{d}$, the KL-SBM increases with the square of the displacement $d^{2}$, demonstrating a sensitive measure of even minor positional changes.
The $\mathscr{S}^{JS}_{T_{\bm{d}}}[\mu]$ case will be more challenging because it involves a Gaussian mixture $m=\frac{1}{2}\mu+\frac{1}{2}T_{\bm{d}}[\mu]$. We compute this numerically and show the result in Figure 1(a) for various values of $U_{\text{iso}}$. The JS-SBM increases approximately quadratically with increasing $\bm{d}$. As we might expect, it goes up more slowly when $U_{\text{iso}}$ is larger. In Figure 1(b), we consider a threshold value of $\bm{d}$ at which the JS-SBM reaches $0.1$ and plot this as a function of $U_{\text{iso}}$ $=U$. The slope of the plot indicates that achieving a JS-SBM of $0.1$ requires an approximate increase of $0.03$ in the displacement magnitude $d$ for every $0.01$2 increase in $U$.
@@ 4 Numerical Methods
$\begin{split}\mathscr{S}_{T_{\alpha}}^{KL}[\mu]&=\int_{\mathbb{R}^{3}}\mu(x)\log\frac{\mu(x)}{T_{\alpha}[\mu](x)}\,dx\\ &=\mathbb{E}_{x\sim\mu}\left[\log\frac{\mu(x)}{T_{\alpha}[\mu](x)}\right].\end{split}$ (16)
@@
1. Draw $M$ independent and identically distributed (i.i.d.) samples $x_{i}$ from the probability density function $\mu$.
2. Compute $\left(\mathscr{S}_{T_{\alpha}}^{KL}\right)^{MC}[\mu]=\frac{1}{M}\sum_{i=1}^{M}h(x_{i})$, where $h(x_{i})=\left[\log\left({\frac{\mu(x_{i})}{T_{\alpha}[\mu](x_{i})}}\right)\right]$.
@@
1. Choose an arbitrary large $M^{\prime}$, and draw $M^{\prime}$ i.i.d. samples $x_{i}$ from $\mu$.
2. Compute $\sigma^{2}=\text{Var}(h(x_{i}))$, where $h(x_{i})=\left[\log\left({\frac{\mu(x_{i})}{T_{\alpha}[\mu](x_{i})}}\right)\right]$.
@@
5 Numerical Results
This section discusses how the SBM quantitatively analyzes distortions in finite clusters. When focusing on local distortions, we often examine smaller, locally distorted segments within larger structures. Here, we present two finite clusters as examples, each a segment cut from infinite crystal structures, to demonstrate the practical use of SBM for local distortion analysis. The first cluster, discussed in Section 5.1, is taken from an FCC Nickel crystal structure. Its high inherent symmetry makes it ideal for evaluating how SBM detects symmetry violations. We also examine how different structural cutout choices introduce SB, as described in Section 5.1.1. The second example involves a supercell derived from a distorted perovskite with octahedral tilts, detailed in Section 5.2. This case illustrates how SBM enables dynamic analysis of distortion processes by tracking SBM as the perovskite transitions from its original, undistorted state to one with octahedral tilts.
5.1 The violation of symmetry elements of local distorted Nickel
5.1.1 The symmetry breaking from the boundary of the finite cluster

### v2

@@ I Introduction
Our statistical symmetry breaking measure (SBM), $\mathscr{S}_{T_{\alpha}}$, uses information theory to quantify symmetry loss in a finite atomic cluster under a transformation $T_{\alpha}$. The cluster is represented by a normalized electron-weighted density function $\mu$. The measure is defined as the Jensen-Shannon (JS) divergence [20] between the transformed density $(T_{\alpha})_{\#}\mu$ and the original $\mu$, providing a quantitative assessment of the structure’s deviation from the symmetry element $T_{\alpha}$. The measure can be computed for any operation in any atomic cluster. For an undistorted crystal structure, it returns zero for each crystallographic symmetry operation by design. However, with finite size clusters whose shape breaks symmetry, unit cell distortions, or atomic displacements, it yields a positive value that indicates the degree of SB with respect to each symmetry operation.
In this paper, we derive this measure and conduct numerical tests to evaluate its behavior for different cases, comparing it with another related SBM based on the Kullback-Leibler (KL). We demonstrate that these measures are useful for investigating factors affecting SB, such as boundary shape, perturbation size, and distorted atoms. They also help identify which symmetries are most violated or preserved when a cluster experience a specific distortion.
II KL and JS divergences in symmetry breaking
@@
$\mathscr{S}_{T_{\alpha}}=\int_{D}\mu(\zeta)\ln\left(\frac{\mu(\zeta)}{(T_{\alpha})_{\#}\mu(\zeta)}\right)\,d\zeta,$ (1)
@@
$\begin{split}D_{\text{JS}}(P,Q)=&\frac{1}{2}D_{\text{KL}}(P\parallel M)+\frac{1}{2}D_{\text{KL}}(Q\parallel M)\\ =&\frac{1}{2}\int_{\mathcal{X}}p(x)\log\frac{p(x)}{m(x)}\,dx+\frac{1}{2}\int_{\mathcal{X}}q(x)\log\frac{q(x)}{m(x)}\,dx,\end{split}$ (3)
@@
III A symmetry breaking measure for finite clusters
III.1 Finite cluster representation
In our proposed model, we represent a finite cluster of atoms as a normalized electron-weighted atomic density function $\mu(\bm{x}):\bm{x}\in\mathbb{R}^{3}\to\mathbb{R}$. The electron density at each atom is assumed to be located at the position of the atomic nucleus. However, due to atomic motions, the probability distribution of the atomic density can be approximated by a three-dimensional Gaussian distribution, commonly known as the Debye-Waller approximation [31], weighted by the number of electrons held by the atom. Specifically, for any atom $k$ where the average location of the nucleus is at $\bm{x}_{k}$, the electron-weighted atomic density can be written as
@@ III.1 Finite cluster representation
where, $e_{k}$ denotes the number of electrons held by atom $k$, $o_{k}$ is its occupancy factor, and $\bm{U}_{k}\in\mathbb{R}^{3\times 3}$ is the anisotropic atomic displacement tensor (ADT). For simplicity, here we treat this distribution as being isotropic, that is, $\bm{U}_{k}=U_{k}\cdot\bm{I}_{3}$ is a diagonal matrix, and all the diagonal entries, $U_{k}$, are equal to the $U_{\text{iso}}$ of atom $k$. Here $U_{k}=1/3\times(\bm{U}_{k,11}+\bm{U}_{k,22}+\bm{U}_{k,33})$ can be interpreted as a mean-square displacement averaged over all the three directions [27].
@@
$\tilde{\mu}=\sum_{k=1}^{N}\tilde{\mu_{k}}=\sum_{k=1}^{N}\tilde{\phi_{k}}\,\mathcal{N}(\bm{x}_{k},U_{k}\cdot\bm{I}_{3}),$ (5)
@@
$\mu=\frac{\tilde{\mu}}{\|\tilde{\mu}\|_{L_{1}}}=\frac{\tilde{\mu}}{\sum_{k=1}^{N}\tilde{\phi_{k}}}=\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(\bm{x}_{k},U_{k}\cdot\bm{I}_{3}),$ (6)
@@
$\begin{split}(T_{\alpha})_{\#}\mu=&(T_{\alpha})_{\#}\left(\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(\bm{x}_{k},U_{k}\cdot\bm{I}_{3})\right)\\ =&\sum_{k=1}^{N}\phi_{k}\,\mathcal{N}(T_{\alpha}(\bm{x}_{k}),U_{k}\cdot\bm{I}_{3}).\end{split}$ (7)
III.2 Symmetry breaking measure
@@
$\begin{split}\mathscr{S}_{T_{\alpha}}^{KL}[\mu]=&D_{\text{KL}}\left(\mu\|(T_{\alpha})_{\#}\mu\right)\\ =&\int_{\mathbb{R}^{3}}\mu(\bm{x})\log\left({\frac{\mu(\bm{x})}{(T_{\alpha})_{\#}\mu(\bm{x})}}\right)\,d\bm{x}.\end{split}$ (8)
@@
$\mathscr{S}_{T_{\alpha}}^{JS}[\mu]=\frac{1}{2}\int_{\mathcal{X}}\mu(\bm{x})\log\frac{\mu(\bm{x})}{m(\bm{x})}\,dx+\frac{1}{2}\int_{\mathcal{X}}(T_{\alpha})_{\#}\mu(\bm{x})\log\frac{(T_{\alpha})_{\#}\mu(\bm{x})}{m(\bm{x})}\,d\bm{x},$ (9)
where $m$ is an equal mixture of $\mu$ and $(T_{\alpha})_{\#}\mu$. Specifically,
$m=\sum_{k=1}^{N}\frac{\phi_{k}}{2}\,\left[\mathcal{N}(\bm{x}_{k},U_{k}\cdot\bm{I}_{3})+\mathcal{N}(T_{\alpha}(\bm{x}_{k}),U_{k}\cdot\bm{I}_{3})\right].$ (10)
Although the pointwise log‑ratio can be negative, the integral is equal to a JS-divergence between $\mu$ and $(T_{\alpha})_{\#}\mu$ and is therefore non‑negative by Gibbs’ inequality [20]. $\mathscr{S}_{T_{\alpha}}^{KL}[\mu]$ and $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ both satisfy the three divergence properties:
@@
  • Self identification: $\mathscr{S}_{T_{\alpha}}[\mu]=0$ only if $(T_{\alpha})_{\#}\mu=\mu$.
@@
For the case of the JS-SBM, when using a logarithm with a base of 2 in Eq. 9, it is bounded by one. If the natural logarithm is employed, the upper bound becomes $\sqrt{2}$. $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ reaches its upper bound if and only if $\mu$ and $T_{\alpha}$ are disjoint, indicating that the operation $T_{\alpha}$ is broken completely. For a Gaussian mixture $\mu$, which is consistently non-zero, this upper bound will not be realized. However, if $\bm{x}_{k}$ and all $T_{\alpha}(\bm{x}_{k})$ are sufficiently distant in terms of their $U_{\text{iso}}$’s, $\mathscr{S}_{T_{\alpha}}^{JS}[\mu]$ can approach close to its bound.
@@
For a high-symmetry heterostructure, such as an infinite crystal with a Face-Centered Cubic (FCC) lattice of nickel, we refer to a set of symmetry operations denoted as $\mathscr{T}=\{T_{i}\}$, which contains transformations that, when applied, yield an identical object. We then introduce some changes to this heterostructure. This results in the object losing some of the symmetry operations from the set $\mathscr{T}$ (and in principle, possibly gaining some new ones) and having a new set of symmetry operations, $\mathscr{T}^{\prime}=\{T^{\prime}_{i}\}$. Strictly speaking, SB of this structure can be defined as the condition where the SBM $\mathscr{S}_{T_{i}}$ of some symmetry operation $T_{i}\in\mathscr{T}$ result in $\mathscr{S}_{T_{i}}>0$.
III.2.1 Symmetry breaking measure of symmetry operators
@@
Determining the SBM of symmetry operators is not trivial, because the symmetry operators can generate more than one symmetry operation. For example, such as the proper axes of rotation $C_{n}$ can lead to n1n-1 operations, specifically $C_{n},C_{n}^{2},\dots,C_{n}^{n-1}$. $C_{n}^{n}$ is considered the “identity” operation, denoted $E$. For example, in structures with n-fold rotational symmetry $C_{n}$, the SBM for these operations are zero ($\mathscr{S}_{C_{n}}=\mathscr{S}_{C_{n}^{2}}=\dots=\mathscr{S}_{C_{n}^{n-1}}=0$). However, if a perturbation disrupts this n-fold symmetry, the SBM for $C_{n}^{m}$ and $C_{n}^{k}$ may differ for some $m,k\in\{1,\dots,n-1\}$. Therefore, we need a measure to quantify how much a particular symmetry operator, such as n-fold rotation, is broken by a symmetry lowering distortion.
To address this, we introduce symmetry breaking measure of operators, $\mathscr{S}_{\mathscr{O}}$, as the average SBM of all the $|\{\mathscr{O}^{m}\}|$ operations derived from $\mathscr{O}$.
$\mathscr{S}_{\mathscr{O}}=\frac{1}{|\{\mathscr{O}^{m}\}|}\sum_{\mathscr{O}^{m}}\mathscr{S}_{\mathscr{O}^{m}}.$ (12)
For example $\mathscr{S}_{C_{4}}=1/3\sum_{m=1}^{3}\mathscr{S}_{C_{4}^{m}}$. Here, a symmetry operator $\mathscr{O}$ includes its geometric element (e.g., for $C_{n}$ the rotation axis is fixed; for a mirror the plane is fixed). Thus, Eqn (12) averages over the derived operations $\{\mathscr{O}^{m}\}$ of this single fixed element (e.g., $C_{n}^{m}$ about the same axis) and does not average across different axis placements or mirror–plane altitudes. This measure gives a value of 0 when symmetry is preserved, but ranges between 0 and $1$ for the JS-SBM if symmetry is partially broken.
III.2.2 Symmetry breaking measure of a single-atom system
To gain insights, we will consider a structure with a single-atom in the unit cell with $_{\text{iso}} =U=U$ and which becomes displaced by $\bm{d}$. WLOG, let
$\mu=$ $\mathcal{N}(\bm{0},U\cdot\bm{I}_{3}),\,$ (13)
$(T_{\bm{d}})_{\#}\mu=$ $\mathcal{N}(\bm{d},U\cdot\bm{I}_{3}).$ (14)
@@ III.2.2 Symmetry breaking measure of a single-atom system
$\mathscr{S}^{KL}_{T_{\bm{d}}}=D_{\text{KL}}\left(\mu\|(T_{\bm{d}})_{\#}\mu\right)=\frac{d^{2}}{2U},$ (15)
where $d=\|\bm{d}\|_{L_{2}}$ is the displacement distance. Details of the derivation are provided in Appendix A. This result suggests that for small distortions $\bm{d}$, the KL-SBM increases with the square of the displacement $d^{2}$, demonstrating a sensitive measure of even minor positional changes.
The $\mathscr{S}^{JS}_{T_{\bm{d}}}[\mu]$ case will be more challenging because it involves a Gaussian mixture $m=\frac{1}{2}\mu+\frac{1}{2}(T_{\bm{d}})_{\#}\mu$. We compute this numerically and show the result in Figure 1(a) for various values of $U_{\text{iso}}$. The JS-SBM increases approximately quadratically with increasing $d$. As we might expect, it goes up more slowly when $U_{\text{iso}}$ is larger. Panel (a) focuses on the small-$d$ regime to display $d^{*}$ clearly; for completeness, under logarithm base 2, the JS-SBM curves approach an asymptote $\leq 1$ as $d\to\infty$. In Figure 1(b), we consider a threshold value of $d^{*}$ at which the JS-SBM reaches $0.1$ and plot this as a function of $_{\text{iso}} =U=U$. The slope of the plot indicates that achieving a JS-SBM of $0.1$ requires an approximate increase of $0.03$ Å in the displacement magnitude $d$ for every $0.01$ Å2 increase in $U$.
@@ IV Numerical Methods
$\begin{split}\mathscr{S}_{T_{\alpha}}^{KL}[\mu]&=\int_{\mathbb{R}^{3}}\mu(\bm{x})\log\frac{\mu(\bm{x})}{(T_{\alpha})_{\#}\mu(\bm{x})}\,d\bm{x}\\ &=\mathbb{E}_{\bm{x}\sim\mu}\left[\log\frac{\mu(\bm{x})}{(T_{\alpha})_{\#}\mu(\bm{x})}\right].\end{split}$ (16)
@@
1. Draw $M$ independent and identically distributed (i.i.d.) samples $\bm{x}_{i}$ from $\mu$.
2. Compute $\left(\mathscr{S}_{T_{\alpha}}^{KL}\right)^{MC}[\mu]=\frac{1}{M}\sum_{i=1}^{M}h(\bm{x}_{i})$, where $h(\bm{x}_{i})=\left[\log\left({\frac{\mu(\bm{x}_{i})}{(T_{\alpha})_{\#}\mu(\bm{x}_{i})}}\right)\right]$.
@@
1. Choose an arbitrary large $M^{\prime}$, and draw $M^{\prime}$ i.i.d. samples $\bm{x}_{i}$ from $\mu$.
2. Compute $\sigma^{2}=\text{Var}(h(\bm{x}_{i}))$, where $h(\bm{x}_{i})=\left[\log\left({\frac{\mu(\bm{x}_{i})}{(T_{\alpha})_{\#}\mu(\bm{x}_{i})}}\right)\right]$.
@@
V Numerical Results
This section discusses how the SBM quantitatively analyzes distortions in finite clusters. When focusing on local distortions, we often examine smaller, locally distorted segments within larger structures. Here, we present two finite clusters as examples, each a segment cut from infinite crystal structures, to demonstrate the practical use of SBM for local distortion analysis. The first cluster is taken from an FCC Nickel crystal structure. Its high inherent symmetry makes it ideal for evaluating how SBM detects symmetry violations. We also examine how different structural cutout choices introduce SB. The second example involves a supercell derived from a distorted perovskite with octahedral tilts. This case illustrates how SBM enables dynamic analysis of distortion processes by tracking SBM as the perovskite transitions from its original, undistorted state to one with octahedral tilts.
V.1 The violation of symmetry elements of local distorted Nickel
V.1.1 The symmetry breaking from the boundary of the finite cluster